OpenStudy (kmazraee):

I want the Lorentz Transformation of the Maxwell's Equation. anyone have a link or PDF file or describing here?

OpenStudy (kmazraee):

No! I want it with simple mathmatics, not as Tensors!

OpenStudy (vincent-lyon.fr):

Maxwell's equations are unchanged by a Lorentz transformation: laws of electromagnetism are the same in all inertial frames.

OpenStudy (vincent-lyon.fr):

If what you want to know is how Lorentz transformation affects E and B fields, it goes: \(E'_x = E_x\) \(E'_y = \gamma \:( E_y - c \beta B_z ) \) \(E'_z = \gamma \:( E_z + c \beta B_y ) \) \(B'_x = B_x \) \(B'_y = \gamma \:( B_y + \frac{\beta}{c} E_z )\) \(B'_z = \gamma \:( B_z - \frac{\beta}{c} E_y ) \)

OpenStudy (kmazraee):

@Vincent-Lyon.Fr can you describe me in a simple mathematical structure?(no tensor uses) please solve it :D

OpenStudy (unklerhaukus):

\[\dagger\]

OpenStudy (anonymous):

http://www.fourmilab.ch/etexts/einstein/specrel/www/ Einstein's paper

OpenStudy (anonymous):

Vincent gave you the transformation by components. What more do you want?